OpenStudy (anonymous):

check that if alpha is a real solution to the quadratic eq. x^2+bx+c = 0 , then y=e^(alpha*t) is a solution to the ordinary diff. eq. y''+by'+cy = 0

OpenStudy (anonymous):

yeah, that's called characteristic equation

OpenStudy (anonymous):

that doesn't really help me at all

OpenStudy (anonymous):

you want to prove it right?

OpenStudy (anonymous):

yeah but we haven't learned anything about characteristic equations

OpenStudy (anonymous):

well , if you have differential equation of form a y''+ b y' +c y=0 you can write it as algebraic equation and solve a x^2 + b x^2 +c =0

OpenStudy (anonymous):

why do we do that? well , we assume soloution would be in form y= e^at so y'=a e^at y''= a^2 e^(at) r y'' +s y'+t y =0 r (a^2 e^(at))+ s(a e^(at))+t(e^at)=0 divide by e^(a t) r a^2 + s a+ t=0

OpenStudy (anonymous):

instead of dividing both sides by e^(at) could you factor it out then realize that you end up with something that looks like the quadratic equation given in the earlier equation and since that =0 then multiplying e^(at) by 0 = 0

OpenStudy (anonymous):

yes, that's better

OpenStudy (anonymous):

thanks!